the sum of the two numbers is -3, while their product - 56. What are those numbers
thank you

Answer by Boreal(15235)

x+y=-3
xy=-56
so x=-56/y
and y-(56/y)=-3
so y^2-56=-3y
and y^2+3y-56=0
y=(1/2)(-3+/- sqrt (233))=(1/2)(-3+/-15.26)=-9.13 and 6.13 to two decimal places.